OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4.

OpenStudy (nurali):

f(x) = 2x + 2 assuming you mean inverse: f-1(x) = (x - 2)/2 f-1(4) = (4-2)/2 = 1

OpenStudy (anonymous):

1